OpenStudy (anonymous):

You earn 245$ in gross pay a week and 2/5 of your check is taken out for taxes and deductions. How much of your check is taken out for taxes and deductions

OpenStudy (texaschic101):

2/5 of 245 = 2/5 x 245/1 = 98 $ 98 dollars of your check is taken out leaving you with $ 147 check... 2/5(245) + 3/5(245) = 245 98 + 147 = 245 245 = 245 (correct)